OpenStudy (anonymous):

A certain waterfall is 145.2 m high and has water flow rate of 39 m3/s. The acceleration of gravity is 9.8 m/s2 . Find the maximum electric power that can be generated by these falls assuming 100% conversion of mechanical energy to electric energy. (Take the density of water to be 1.00 × 103 kg/m3) Answer in units of W.

OpenStudy (anonymous):

pot energy =mgh = now density * flow rate =1000*39 =39000 Kg/s Energy =pot/time=mgh/t=39000*145.2*9.8=55495440 W